OpenStudy (anonymous):

Solve for R in the formula V = IR, when I = 5 and V = 15.

OpenStudy (anonymous):

R=3

OpenStudy (anonymous):

you do v divided by i to get r

OpenStudy (solomonzelman):

You just lug the numbers into the formula V = I × R ♂ ♂ ♂ 15 5 ? 15=5 × R (R is an unknown number) divide each side by 5, (i.e. divide 15 by 5 and divide (5 × R) by 5... you get?

OpenStudy (solomonzelman):

I meant to say plug (in the beginning), not lug.

OpenStudy (anonymous):

yeah @Deaner 15/5 is 3

OpenStudy (anonymous):

yea i was just telling him how you did it